OpenStudy (anonymous):

7(4 -3z) + 3 = -6(z - 4) - 8

OpenStudy (anonymous):

if I am correct... this equation can not be solved

OpenStudy (raffle_snaffle):

it can be solved

OpenStudy (raffle_snaffle):

distribute the 7 throug and -6 through then combine like terms

OpenStudy (anonymous):

I know the next step is 28 - 21z + 3 = -6z - -24 - 8

OpenStudy (paki):

next step is combine like terms... like "z" here....

OpenStudy (paki):

like this... -21z+6z=-24-8-28-3

OpenStudy (paki):

now solve... what you get...

OpenStudy (anonymous):

but the 6z is a negative

OpenStudy (paki):

yup 6z is negative, but when you shift it to the other side, its sign changes...

OpenStudy (paki):

-15z=-63.... is this rite.... @patrick_alvarado89

OpenStudy (anonymous):

yea

OpenStudy (paki):

hmm... so what will be the next step....?

OpenStudy (anonymous):

-15z/-15z = -63/-15

OpenStudy (paki):

hmmm excellent rite... so answer will be....?

OpenStudy (paki):

divide by only "15",,,, not "z"... ok...

OpenStudy (anonymous):

-21/-5

OpenStudy (paki):

see... this is -15z=-63... ok... so first cancel the "minus" signs from bith thwe sides, then divide by "15"... like this.... 15z=63 or z=63/15 or z=21/5

OpenStudy (paki):

clear now... @patrick_alvarado89

OpenStudy (anonymous):

ok... I thought that since they were negatives they stayed negatives

OpenStudy (paki):

no... negative signs will be cancel, as they are on both the sides... if there were only one negative sign, then it will be as it as...

OpenStudy (anonymous):

got it.

OpenStudy (paki):

nice.... :-)

OpenStudy (anonymous):

thanks. I owe you

OpenStudy (paki):

:)